INFRARED ACTION SPECTROSCOPY AND TIME-RESOLVED DYNAMICS OF THE OD-CO REACTANT COMPLEX

\underline{ILANA B. POLLACK}, MARIA TSIOURIS, HELEN O. LEUNG,^{a} AND MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

Infrared action spectra of the linear OD-CO reactant complex have been recorded in the OD overtone region near 1.9 μm using an infrared pump-ultraviolet probe technique. The pure overtone band of OD-CO (2ν_{OD}) was observed at 5148.4 cm^{−1}, and combination bands involving the simultaneous excitation of OD stretch and D-atom bend were identified 200.7 and 232.1 cm^{−1} to higher energy. Band assignments and spectroscopic constants have been derived from the rotationally resolved structure of the spectra. Direct time-domain measurements yielded a lifetime of 34 ns for OD-CO (2ν_{OD}) prior to decay via inelastic scattering or chemical reaction. This is significantly longer than the laser-limited lifetime of ≤ 5 ns observed for OH-CO (2ν_{OH}), and is attributed to the closing of a near resonant vibrational energy transfer channel upon deuteration. Intermolecular bending excitation, which drives the structural transformation from the reactant complex to the transition state for reaction, results in a dramatic shortening of the lifetime to ≤ 5 ns. Excitation of the D-atom bend also supplies sufficient energy to reopen the near resonant vibrational energy transfer channel. Finally, an OD-CO binding energy of D_{0} ≤ 500 cm^{−1} has been established from the OD (v=1) product state distribution observed following infrared overtone excitation.

aOn sabbatical leave from Department of Chemistry, Mount Holyoke College, South Hadley, Massachusetts 01075-6407